# Towards Robust Infrared Small Target Detection: A Feature-Enhanced and Sensitivity-Tunable Framework

### v2

@@ 4.4 Effect verification of the multi-scale fusion
To evaluate the effectiveness of the multi-scale fusion strategy, we conduct extensive experiments on the W-IRSTD dataset, using MSDA-Net as the basic model and EEDM Loss as the loss function. Specifically, we resize the original images to various resolutions, including 256×256, 384×384, 512×512, 640×640, 768×768, 896×896, and 1024×1024, to systematically explore the impact of input scale on model performance. Based on this, we take the best-performing single-scale model as the foundation and progressively integrate suboptimal models to further evaluate the effectiveness of the multi-scale fusion strategy. For clarity, each model is named based on its corresponding input resolution. For example, the model trained with 256×256 images is denoted as the 256 model.
@@
From table 2, the 1024 model achieves the best performance. Compared with the worst performing 256 model, the 1024 model improves the $IoU$, $P_{d}$, and Score by 37.35% (from 43.78 to 60.13), 56.25% (from 48.75 to 76.17) and 47.29% (from 46.27 to 68.15), respectively. The analysis reveals that when images from the W-IRSTD dataset are downsampled to smaller resolutions, such as 256×256, the model performance drops significantly. This degradation is primarily attributed to the loss of pixel-level information during the downsampling process, which causes small targets to shrink or even vanish entirely, making it difficult for the network to extract effective features. In addition, considering that the proposed AS strategy will be used later, it can effectively improve the detection rate of the model. Therefore, to ensure the segmentation accuracy, we choose the 896 model with the best $IoU$ metrics as the basis model, and fuse the suboptimal scale models in the $IoU$ metrics one by one. The experimental results are shown in table 5.
@@
From table 5, the fusion of the 896, 768, and 1024 models yields the most significant performance improvement. Specifically, compared with the 896 model, which has the best performance at a single scale, the fusion of the 896 model, the 768 model and the 1024 model can improve its $IoU$, $P_{d}$ and Score by 4.20% (from 61.89 to 64.49), 0.88% (from 71.95 to 72.58) and 2.42% (from 66.92 to 68.54), respectively. At the same time, we find that when the performance of the fused scale models degrades severely, it will lead to the degradation of the performance of the final model. In addition, the multi-scale fusion model will perform better on $F_{a}$, which shows that the multi-scale fusion strategy is conducive to reducing the missed detections and further improving the reliability of the model. To more intuitively demonstrate the effect of multi-scale model fusion, we show the prediction results of the 896 model, 768 model, 1024 model and their fusion model in fig. 6. It is worth noting that although multi-scale model fusion significant improvements in detection performance, it inevitably increases inference time. Therefore, a reasonable trade-off between accuracy and efficiency should be made based on specific requirements in practical applications to achieve optimal overall performance.
@@
Figure 7: 2D comparison of the effects of the AS strategy. In each group of images, the ones with green borders represent detection results without the AS strategy, while the ones with red borders represent detection results after applying the AS strategy. Local zoom has been applied to emphasize the differences in key areas.
4.5 Effect verification of the AS strategy
To verify the AS strategy, we adjust the thresholds th1 and th2 on the optimal multi-model fusion scheme (“896+768+1024”). Table 5 shows the experimental results obtained by adjusting the threshold th1 alone without adjusting the threshold th2. Table 5 shows the experimental results obtained by adjusting the threshold th2 on the basis that th1 is set to the optimal value in table 5.
From table 5, when the threshold th1 is set to 0.3, the model achieves the best overall performance. Compared to the conventional setting of th1 at 0.5, adjusting th1 to 0.3 leads to 5.43% (from 68.54 to 72.26) improvement in the comprehensive metric Score. The key to this improvement lies in the model’s ability to significantly enhance the detection rate $P_{d}$ without notably affecting segmentation accuracy. Specifically, $P_{d}$ increases by 10.62% (from 72.58 to 80.29). Notably, when the threshold th1 is reduced to 0.05, $F_{a}$ will surge and exceed the constraint range.
From table 5, the introduction of th2 significantly improves the detection rate of the network while maintaining segmentation accuracy for infrared small targets and effectively controlling the false alarm rate. On the W-IRSTD dataset, based on the multi-scale fusion scheme of “896+768+1024”, the model achieves optimal performance when th1 is set to 0.3 and th2 is set to 0.1. Specifically, compared to using only th1 set to 0.3, the further introduction of th2 set to 0.1 leads to a 4.76% (from 72.26 to 75.70) improvement in Score metrics.
To more intuitively demonstrate the effectiveness of the AS strategy, we present some results as shown in fig. 7. For extended targets, the appropriate application of the AS strategy helps refine the shape and details of the target. For point or spot targets, the strategy helps improve detection rates and effectively capture weak target features.
Table 6: $IoU$ (%), $P_{d}$ (%) and $F_{a}$$10^{-6}$) values of different methods achieved on the separate NUAA-SIRST, NUDT-SIRST, and IRSTD-1k datasets. (213:214), (663:664) and (800:201) denote the division of training samples and test samples. Red denotes that the $F_{a}$ exceeds the acceptable upper limit, making the results for this group invalid. Green denotes the optimal result.
@@ 4.5 Effect verification of the AS strategy
Scheme Description $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score
Original 62.24 94.30 111.84 - 65.36 96.08 37.48 80.72 61.08 93.60 22.51 77.34
@@
MSDA-Net [32] +FEST 80.51 97.72 15.23 89.12 94.39 99.68 3.13 97.04 71.19 97.98 25.07 84.59
@@
Figure 8: Visualization of several SIRST methods on the NUAA-SIRST dataset. Red denotes the correct detections, blue denotes the false detections, and yellow denotes the missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
4.6 Validation of the FEST Framework on multiple datasets
To validate the effectiveness and robustness of the proposed FEST framework, we integrate multiple advanced SIRST networks into the framework and conduct experiments on three public datasets. To ensure a fair comparison, all models are retrained under the same settings, with the number of training epochs uniformly set to 300 to reduce training costs. This setup differs from that used in our previous work. In addition, we explore the performance gain brought by using only EEDM Loss and AS strategy in the FEST framework, with almost no additional resource consumption. This lightweight variant is referred to as the LW-FEST framework. The thresholds th1 and th2 in both the FEST and LW-FEST frameworks are set to 0.5 and 0.1, respectively. Notably, considering the characteristics of each dataset, we adopt differentiated configurations in the multi-scale fusion strategy: for NUAA-SIRST and IRSTD-1k, we fuse models with sizes of 512 and 768, whereas for NUDT-SIRST, we fuse models with sizes of 256 and 384.
@@ 4.6 Validation of the FEST Framework on multiple datasets
Figure 9: Visualization of several SIRST methods on the NUDT-SIRST dataset. Red denotes the correct detections, blue denotes the false detections, and yellow denotes the missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
@@
Figure 10: Visualization of several SIRST methods on the IRSTD-1k dataset. Red denotes the correct detections, blue denotes the false detections, and yellow denotes the missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
From table 6, embedding various SIRST networks into the lightweight LW-FEST framework can effectively improve the model performance without increasing resource consumption. These results demonstrate that the LW-FEST framework exhibits strong robustness and generalization capabilities. It can be seamlessly integrated with various SIRST networks, effectively enhancing their feature extraction and representation abilities. Meanwhile, when networks with relatively poor performance (such as ACM) are embedded into the framework, the performance improvement is more significant. In addition, the application of the complete FEST framework can further improve the detection performance. Specifically, on the NUAA-SIRST dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 0.46% - 6.89% in the evaluation metric Score. On the NUDT-SIRST dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 0.48% - 6.34% in the evaluation metric Score. On the IRSTD-1k dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 1.66% - 5.56% in the evaluation metric Score.
To further demonstrate the superiority of the proposed framework, we present visual comparisons of partial difficult samples detection results across three datasets, as shown in fig. 8, fig. 10, and fig. 10. Equipping SIRST networks with the FEST framework leads to clearer target edges, more comprehensive detection results, and enhanced perception and localization accuracy for small targets in diverse detection scenarios.

### v3

@@ 4.4 Effect verification of the multi-scale fusion
To evaluate the effectiveness of the multi-scale fusion strategy, we conduct extensive experiments on the W-IRSTD dataset, using MSDA-Net as the basic model and EEDM loss as the loss function. Specifically, we resize the original images to various resolutions, including 256×256, 384×384, 512×512, 640×640, 768×768, 896×896, and 1024×1024, to systematically explore the impact of the input scale on model performance. On this basis, we take the best-performing single-scale model as the foundation and progressively integrate suboptimal models to further evaluate the effectiveness of the multi-scale fusion strategy. For clarity, each model is named based on its corresponding input resolution. For example, the model trained with 256×256 images is denoted as the 256 model.
@@
From Table 2, the 1024 model achieves the best performance. Compared with the worst performing 256 model, the 1024 model improves the $IoU$, $P_{d}$, and Score by 37.35% (from 43.78 to 60.13), 56.25% (from 48.75 to 76.17) and 47.29% (from 46.27 to 68.15), respectively. The analysis reveals that when images from the W-IRSTD dataset are downsampled to smaller resolutions, such as 256×256, the model performance drops significantly. This degradation is primarily attributed to the loss of pixel-level information during the downsampling process, which causes small targets to shrink or even vanish entirely, making it difficult for the network to extract effective features. In addition, considering that the proposed AS strategy will be used later, it can effectively improve the detection rate of the model. Therefore, to ensure segmentation accuracy, we choose the 896 model with the best $IoU$ metric as the basis model, and fuse the suboptimal scale models in the $IoU$ metric one by one. The experimental results are shown in Table 5.
@@
From Table 5, the fusion of the 896, 768, and 1024 models yields the most significant performance improvement. Specifically, compared with the 896 model, which has the best performance at a single scale, the fusion of the 896 model, the 768 model and the 1024 model can improve its $IoU$, $P_{d}$ and Score by 4.20% (from 61.89 to 64.49), 0.88% (from 71.95 to 72.58) and 2.42% (from 66.92 to 68.54), respectively. At the same time, we find that when the performance of the fused scale models degrades severely, it will lead to the degradation of the performance of the final model. In addition, the multi-scale fusion model will perform better on $F_{a}$, which shows that the multi-scale fusion strategy is conducive to reducing the missed detections and further improving the reliability of the model. To more intuitively demonstrate the effect of multi-scale model fusion, we show the prediction results of the 896 model, 768 model, 1024 model and their fusion model in Fig. 8. It is worth noting that although multi-scale model fusion achieves significant improvements in detection performance, it inevitably increases the inference time. Therefore, a reasonable trade-off between accuracy and efficiency should be made based on specific requirements in practical applications to achieve optimal overall performance.
4.5 Effect verification of the AS strategy
To validate the AS strategy, we adjust the thresholds $th1$ and $th2$ on the optimal multi-model fusion scheme (“896+768+1024”). Table 5 shows the experimental results obtained by adjusting the threshold $th1$ alone without adjusting the threshold $th2$. Table 5 shows the experimental results obtained by adjusting the threshold $th2$ on the basis that $th1$ is set to the optimal value in Table 5.
From Table 5, when the threshold $th1$ is set to 0.3, the model achieves the best overall performance. Compared with the conventional setting of $th1$ at 0.5, adjusting $th1$ to 0.3 leads to 5.43% (from 68.54 to 72.26) improvement in the comprehensive metric Score. The key to this improvement lies in the model’s ability to significantly increase the detection rate $P_{d}$ without notably affecting the segmentation accuracy. Specifically, $P_{d}$ increases by 10.62% (from 72.58 to 80.29). Notably, when the threshold $th1$ is reduced to 0.05, $F_{a}$ will surge and exceed the constraint range.
From Table 5, the introduction of $th2$ significantly improves the detection rate of the network while maintaining segmentation accuracy for infrared small targets and effectively controlling the false alarm rate. On the W-IRSTD dataset, the model achieves optimal performance under the multi-scale fusion scheme “896+768+1024” when $th1$ is set to 0.3 and $th2$ is set to 0.1. Specifically, compared with using only $th1$ set to 0.3, the further introduction of $th2$ set to 0.1 leads to a 4.76% (from 72.26 to 75.70) improvement in Score metric.
To more intuitively demonstrate the effectiveness of the AS strategy, we present some results as shown in Fig. 9. For extended targets, the appropriate application of the AS strategy helps refine the shape and details of the target. For point or spot targets, this strategy helps improve detection rates and effectively capture weak target features.
4.6 Validation of the FEST Framework on multiple datasets
To validate the effectiveness and robustness of the proposed FEST framework, we integrate multiple advanced SIRST networks into the framework and conduct experiments on three public datasets. To ensure a fair comparison, all the models are retrained under the same settings, with the number of training epochs uniformly set to 300 to reduce training costs. This setup differs from that used in our previous work. In addition, we explore the performance gain brought by using only the EEDM loss and AS strategy in the FEST framework, with almost no additional resource consumption. This lightweight variant is referred to as the LW-FEST framework. The thresholds $th1$ and $th2$ in both the FEST and LW-FEST frameworks are set to 0.5 and 0.1, respectively. Notably, considering the characteristics of each dataset, we adopt differentiated configurations in the multi-scale fusion strategy: for NUAA-SIRST and IRSTD-1k, we fuse models with sizes of 512 and 768, whereas for NUDT-SIRST, we fuse models with sizes of 256 and 384.
From Table 6, embedding various SIRST networks into the lightweight LW-FEST framework can effectively improve the model performance with minimal additional resource consumption. These results demonstrate that the LW-FEST framework exhibits strong robustness and generalization, and can be seamlessly integrated with various SIRST networks, effectively enhancing their feature extraction and representation abilities. Meanwhile, when networks with relatively poor performance (such as ACM) are embedded into the framework, the performance improvement is more significant. In addition, the application of the complete FEST framework can further improve the detection performance. Specifically, on the NUAA-SIRST dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 0.46% - 7.40% in the evaluation metric Score. On the NUDT-SIRST dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 0.48% - 6.34% in the evaluation metric Score. On the IRSTD-1k dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 1.66% - 7.26% in the evaluation metric Score.
@@ 4.6 Validation of the FEST Framework on multiple datasets
Figure 9: 2D comparison of the effects of the AS strategy. In each group of images, the ones with green borders represent the detection results without the AS strategy, while the ones with red borders represent the detection results after applying the AS strategy. Local zoom has been applied to emphasize the differences in key areas.
To further demonstrate the superiority of the proposed framework, we present visual comparisons of detection results on several samples across three datasets, as shown in Fig. 10, Fig. 12, and Fig. 12. Equipping SIRST networks with the FEST framework leads to clearer target edges, more comprehensive detection results, and enhanced perception and localization accuracy for small targets in diverse detection scenarios.
Table 6: IoUIoU (%), $P_{d}$ (%) and $F_{a}$$10^{-6}$) values of different methods achieved on the separate NUAA-SIRST, NUDT-SIRST, and IRSTD-1k datasets. (213:214), (663:664) and (800:201) denote the division of training samples and test samples. Red denotes that the $F_{a}$ exceeds the acceptable upper limit, making the results for this group invalid. Green denotes the best result within each group, where each group comprises a method’s Original, +LW-FEST, and +FEST variants.
@@
Scheme Description IoUIoU $P_{d}$ $F_{a}$ ScoreScore IoUIoU $P_{d}$ $F_{a}$ ScoreScore IoUIoU $P_{d}$ $F_{a}$ ScoreScore
Original 62.24 94.30 111.84 - 65.36 96.08 37.48 80.72 60.30 89.56 20.38 74.93
@@
MSDA-Net [zhao2024multi] +FEST 80.51 97.72 15.23 89.12 94.39 99.68 3.13 97.04 71.19 97.98 25.07 84.59
@@
Figure 10: Visualization of several SIRST methods on the NUAA-SIRST dataset. Red denotes correct detections, blue denotes false detections, and yellow denotes missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
4.7 Discussion
1) Why can the EEDM loss generally and effectively improve the performance of SIRST network? The proposed EEDM loss is designed for the “small and weak” characteristics of infrared small targets. Its core mechanism consists of two parts: edge pixel enhancement and difficult pixel mining. Specifically, edge pixel enhancement refers to assigning higher weights to the boundary regions of the target, thereby enhancing the model’s sensitivity to target boundaries. Difficult pixel mining refers to discarding a certain proportion of simple samples, so that network training focuses more on difficult areas, thereby prompting the network to learn more discriminative features. The EEDM loss function enables the network to obtain effective supervision signals and achieve effective performance improvement, without requiring any modifications to the network structure or introducing extra overhead during the inference phase.
2) What is the core value of the AS strategy? The AS strategy breaks through the inherent limitations of single threshold segmentation and fully explores and utilizes the information in the probability mask. Specifically, for SIRST detection scenarios, we introduce the novel concepts of strong targets and weak targets, and adopt a dual threshold mechanism to divide the probability mask into high-confidence strong target regions and relatively low-confidence weak target regions. This strategy preserves the fine segmentation of strong targets while recovering potential missed weak targets at very low cost, thereby improving the overall recall rate and detection performance. In addition, this strategy exhibits good generality and can be seamlessly integrated into existing networks.
@@ 4.7 Discussion
Figure 11: Visualization of several SIRST methods on the NUDT-SIRST dataset. Red denotes correct detections, blue denotes false detections, and yellow denotes missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
@@
Figure 12: Visualization of several SIRST methods on the IRSTD-1k dataset. Red denotes correct detections, blue denotes false detections, and yellow denotes missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
Table 7: Performance comparison of the original networks and their LW-FEST and FEST variants on the IRSTD-1k dataset across different difficulty subsets (easy, medium, and hard). $D$ denotes composite difficulty of a sample. NumNum denotes the number of samples in each subset. Green denotes the best result within each group, where each group comprises a method’s Original, +LW-FEST, and +FEST variants.
Easy ($D\leq 0.3$), NumNum=2 Medium ($0.3<D<0.7$), NumNum=102 Hard ($D\geq 0.7$), NumNum=97
Scheme Description IoU $P_{d}$ $F_{a}$ Score IoU $P_{d}$ $F_{a}$ Score IoU $P_{d}$ $F_{a}$ Score
Original 94.44 100.00 0.00 97.22 72.87 95.24 6.28 84.06 65.26 93.92 83.25 79.59
+ LW-FEST 100.00 100.00 0.00 100.00 73.46 96.60 6.17 85.03 66.06 94.59 72.91 80.33
DNA-Net +FEST 100.00 100.00 0.00 100.00 73.48 96.60 4.97 85.04 67.10 95.95 50.65 81.53
Original 94.74 100.00 0.00 97.37 74.77 94.56 5.12 84.67 62.81 90.54 30.32 76.68
+ LW-FEST 90.00 100.00 0.00 95.00 74.13 96.60 4.08 85.37 62.60 93.24 49.36 77.92
GGL-Net +FEST 94.74 100.00 0.00 97.37 74.32 96.60 4.30 85.46 66.25 93.92 49.12 80.09
Original 90.00 100.00 0.00 95.00 73.58 95.24 22.10 84.41 68.36 93.24 42.63 80.80
+ LW-FEST 100.00 100.00 0.00 100.00 75.84 97.28 4.30 86.56 69.45 93.92 63.95 81.69
UIU-Net +FEST 100.00 100.00 0.00 100.00 74.86 97.96 3.18 86.41 71.11 95.95 57.93 83.53
Original 100.00 100.00 0.00 100.00 73.75 96.60 9.01 85.18 68.68 91.21 32.17 79.95
+ LW-FEST 100.00 100.00 0.00 100.00 73.32 97.28 5.83 85.30 68.78 94.59 53.68 81.69
MSDA-Net +FEST 100.00 100.00 0.00 100.00 73.56 97.96 4.00 85.76 69.11 97.97 47.74 83.54
Table 8: Performance comparison of different networks before and after using LW-FEST and FEST frameworks on the IRSTD-1k dataset. (Params: M, Time: ms).
Original Original + LW-FEST Original + FEST
Method Score Params GFLOPs Time Score Params GFLOPs Time Score Params GFLOPs Time
ACM 74.93 0.40 3.24 1.97 79.02 (↑5.46%) 0.40 3.24 2.00 80.37 (↑7.26%) 0.80 10.54 3.97
ALCNet 77.76 0.43 3.05 1.94 79.94 (↑2.80%) 0.43 3.05 1.99 82.08 (↑5.56%) 0.85 9.90 3.91
MLCL-Net 79.59 0.56 49.71 4.20 80.33 (↑0.93%) 0.56 49.71 4.23 80.91 (↑1.66%) 1.12 161.56 8.44
ALCL-Net 80.43 5.67 54.44 4.12 81.72 (↑1.60%) 5.67 54.44 4.14 82.56 (↑2.65%) 11.34 176.93 8.28
DNA-Net 81.65 4.70 114.26 67.03 82.53 (↑1.08%) 4.70 114.26 67.04 83.20 (↑1.90%) 9.39 371.34 133.98
GGL-Net 81.00 8.99 78.85 15.24 81.36 (↑0.44%) 8.99 78.85 15.26 82.66 (↑2.05%) 17.98 256.25 30.53
UIUNet 82.48 50.54 436.01 16.73 83.99 (↑1.83%) 50.54 436.01 16.75 84.92 (↑2.96%) 101.08 1417.02 33.47
MSDA-Net 82.49 4.81 169.80 16.40 83.45 (↑1.16%) 4.81 169.80 16.42 84.69 (↑2.67%) 9.62 551.85 32.83
3) How does the proposed framework perform on samples of different difficulty levels? To objectively evaluate the detection difficulty of each sample, we define a composite difficulty. It comprises three factors: target size, background complexity, and target-background contrast.
First, to quantify the target size, we compute the pixel area from the target mask, take the logarithm of this area as the scale measure, and apply min-max normalization across the dataset. Considering that smaller targets are generally harder to detect, the size difficulty is defined as:
$D_{t}^{size}=1-Norm(\log\left|{{F_{t}}}\right|)$ (13)
where $F_{t}$ denotes the pixel set of target $t$; $\left|{{F_{t}}}\right|$ denotes its area; $Norm(\cdot)$ denotes the normalization operator; and $D_{t}^{size}$ denotes the size difficulty of target $t$.
Second, to quantify background complexity, we construct a background ring region $R_{t}$} around the target mask via morphological dilation. The ring has a width of $w$ pixels, and any other target regions within it are excluded. In $R_{t}$}, we compute gradient energy to quantify edge and texture complexity and local entropy to quantify intensity uncertainty. The gradient energy is defined as:
$G_{t}=\frac{\sum\limits_{i\in R_{t}}\left((\partial_{x}L(i))^{2}+(\partial_{y}L(i))^{2}\right)}{|R_{t}|}$ (14)
where $R_{t}$} denotes the background ring of target $t$; $\left|{{R_{t}}}\right|$ denotes its area; $L(i)\in[0,1]$ denotes the normalized intensity at pixel $i\in{R_{t}}$; ${\partial_{x}}L(i)$ and ${\partial_{y}}L(i)$ denote the first-order spatial derivatives of the intensity field along the $x$ and $y$ directions, respectively; and $G_{t}$ denotes the gradient energy. The local entropy is defined as:
${H_{t}}=-\sum\limits_{k=1}^{{B_{t}}}{{p_{k}}log({p_{k}})}$ (15)
where ${B_{t}}=\sqrt{\left|{{R_{t}}}\right|}$ denotes the number of histogram bins in $R_{t}$}; $p_{k}$} denotes the probability of the $k_{th}$} bin; and $H_{t}$} denotes the local entropy. Taking both measures into account, the background complexity difficulty is defined as:
$D_{t}^{clutter}=max(Norm({G_{t}}),Norm({H_{t}}))$ (16)
where $\max(\cdot)$ denotes the maximum operator; and $D_{t}^{clutter}$ denotes the background complexity difficulty of the target $t$.
Third, to quantify the target-background contrast, we compute the Signal-to-Clutter Ratio (SCR) using the background ring region $R_{t}$}. The formula of SCR is as follows:
$SC{R_{t}}=\frac{{{\mu_{t}}-{\mu_{{R_{t}}}}}}{{{\sigma_{{R_{t}}}}+\varepsilon}}$ (17)
where $\mu_{t}$} denotes the mean intensity over the target region; $\mu_{{R_{t}}}$} and $\sigma_{{R_{t}}}$} denote the mean and standard deviation within $R_{t}$}; and $\varepsilon$ denotes a small positive constant. The contrast difficulty is defined as:
$D_{t}^{contrast}=1-Norm(SC{R_{t}})$ (18)
where $D_{t}^{contrast}$ denotes the contrast difficulty of target $t$.
Finally, we define the target-level composite difficulty that integrates the three factors above to assess a target’s detection difficulty. The formula is as follows:
${D_{t}}=\alpha D_{t}^{size}+\beta D_{t}^{clutter}+\gamma D_{t}^{contrast}$ (19)
where $\alpha$, $\beta$ and $\gamma$ denote weighting coefficients, which are set to 0.3, 0.3, and 0.4 to balance the contributions of the three factors.
In addition, we define the image-level composite difficulty as follows: If an image contains a single target, the image difficulty equals that target’s difficulty; if multiple targets are present, the image difficulty is defined as the maximum difficulty among all the targets in the image.
To thoroughly evaluate the proposed framework across different difficulty levels (easy, medium, and hard), we stratified the IRSTD-1k dataset into three subsets based on the aforementioned image-level composite difficulty metric and conducted evaluations on each subset. The detailed results are shown in Table 7. The experimental results show that the proposed framework achieves good detection performance across all difficulty levels. In particular, the improvement on the hard subset is more significant.
4) What are the respective computational costs of the FEST and LW-FEST frameworks? To comprehensively evaluate the computational costs and application scenarios of the FEST and LW-FEST frameworks, we conducted comparative experiments on the IRSTD-1k dataset. Specifically, these experiments analyze changes in the Parameters, GFLOPs, and per-image inference time, before and after the introduction of the framework. The detailed results are shown in Table 8. The experimental results demonstrate that after introducing the LW-FEST framework, the number of parameters and GFLOPs remain unchanged, the inference time only increases slightly, but the performance is steadily improved. Notably, compared with the original SIRST network, the SIRST network equipped with the LW-FEST framework achieves a performance improvement of 0.44% - 5.46% in the evaluation metric Score. Therefore, the LW-FEST is particularly suitable for application scenarios with high efficiency requirements. At the same time, introducing the accuracy-oriented FEST framework leads to a relatively significant increase in parameters, GFLOPs, and inference time, while bringing more obvious performance improvements. Specifically, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 1.66% - 7.26% in the evaluation metric Score. Therefore, the FEST is more suitable for application scenarios that prioritize accuracy. Overall, LW-FEST enhances efficiency with negligible computational costs, while FEST further pushes through the performance ceiling through multi-scale fusion, but requires greater resource consumption as a trade-off. These complementary characteristics enable FEST and LW-FEST to meet the requirements of different application scenarios.